Determine the wavelength of a 6000-Hz sound wave traveling along an iron rod. Assume the elastic modulus of iron is 100×109 N/m2.
Answer #1

We know that the speed of sound in iron is given by = sqrt(Y/ρ)

where Y is Young's Modulus(elastic modulus) =100*10^9 N/m^2

ρ is density =7870 kg/m^3

so v = sqrt(100x10^9/7870) = 3564.62 m/s

f = 6000Hz

λ = v/f = 3564.62/6000 = 0.594 m